The combustion of ammonia in the presence of excess oxygen yields NO2 and H2O:

4NH3(g)--4NO2(g)6H2O(g)


Determine the mass in grams of NO2 when 43.9g of NH3 was used.



1
Expert's answer
2021-10-20T05:22:52-0400

"4NH_3\to 4NO_2+6H_2O"

moles of ammonia"=\frac{mass}{RMM}=\frac{43.9}{17}=2.582"

mole ration "NH_3:NO_2=1:1"

moles of "NO_2=2.582"

Mass of "NO_2=number \space of \space moles \times RMM=2.582\times (14+16+16)=118.772g"